# Non-Gaussian state preparation and enhancement using weak-value amplification

### v1

@@ Abstract
We introduce a protocol for generating a broad class of non-Gaussian (nG) quantum states via postselected weak measurement techniques. The scheme involves injecting an arbitrary quantum state and a single photon into the signal and idler ports, respectively, of an interference setup that incorporates a third-order nonlinear medium. A nG state is conditionally produced at the signal output, heralded by the detection of a single photon in one of the idler output channels. The protocol exploits a weak cross-Kerr interaction and effective single-photon nonlinearity enhanced by the weak-value amplification. We show that by tuning the weak value of the photon number operator in the idler mode within experimentally feasible parameters, a wide variety of nG states can be generated with high fidelity. As specific examples, we demonstrate the generation of photon-added states, displaced and squeezed number states, and a continuum of intermediate nG states using coherent and squeezed vacuum inputs, respectively. Furthermore, we show that the protocol enables the enhancement of non-Gaussianity and the enlargement of Schrödinger cat (SC) states when ideal SC states are used as the input. Our results provide an alternative route for the conditional generation of tunable nG states, with potential applications in quantum information processing. This approach may also open new avenues for quantum state engineering using postselected weak measurements.
I Introduction
@@
Non-Gaussian (nG) states (Walschaers, 2021; Walschaers et al., 2020), defined as quantum states in the phase space that deviate from a Gaussian distribution, are often characterized by Wigner function negativity and have garnered significant interest owing to their enhanced utility across various domains of quantum physics. In particular, they play a central role in continuous-variable (CV) quantum information processing (Braunstein and van Loock, 2005), including quantum computation (Lloyd and Braunstein, 1999; Gottesman et al., 2001; Menicucci et al., 2006; Miyata et al., 2016; Mari and Eisert, 2012), quantum metrology (Anisimov et al., 2010; Huver et al., 2008), and foundational studies in quantum theory (Jeong et al., 2003; Stobińska et al., 2007). The generation of nG states requires either nG initial states or nG measurements; the latter alone is sufficient to produce non-Gaussianity. Among the existing techniques, nG unitary transformations and conditional measurements are the two most common approaches (Walschaers, 2021). Over the past decades, considerable efforts have focused on preparing nG states using conditional measurements involving nG quantum operations, such as photon addition and/or subtraction applied to Gaussian states in optical systems (Barbieri et al., 2010; Zavatta et al., 2004; Wakui et al., 2007; Neergaard-Nielsen et al., 2006; Wenger et al., 2004; Chen et al., 2024).
@@
Squeezed single-photon and single-photon-added coherent (SPAC) states are among the most representative nG states and play irreplaceable roles in CV quantum information processing. These states can be prepared via nG operations such as single-photon addition and subtraction, applied to coherent and squeezed vacuum (SV) states, respectively. Another paradigmatic example is the Schrödinger cat (SC) state (Ourjoumtsev et al., 2006) inspired by Schrödinger’s famous thought experiment (Schrödinger, 1935). Current methods for generating SC states include photon-number-resolving detection (Dakna et al., 1997, 1999) and weak and strong nonlinear optical interactions (Yurke and Stoler, 1986; Gerry, 1999). Large-amplitude SC states characterized by coherent amplitudes $|\alpha|\geq 2$, where $\alpha$ is the complex amplitude, are particularly valuable for probing quantum foundations and enabling quantum information protocols (van Enk and Hirota, 2001; Ba An, 2003; An, 2004). However, SC states generated via photon subtraction or third-order optical nonlinearities typically have small amplitudes and fall short of the requirements imposed by advanced quantum tasks.
@@
From a practical perspective, the purity of a state is critical for effective implementation of many quantum information protocols. However, due to limitations in photon detection efficiency and photon-number-resolving capabilities, generating high-fidelity nG states remains significantly more challenging than producing their Gaussian counterparts. This difficulty underscores the limitations of existing methods for preparing high-quality nG states. To address this challenge, one can either enhance the nGity of a given state through optimization techniques or develop alternative approaches capable of generating nG states with high purity and fidelity tailored for specific applications.
@@
In this study, we propose an alternative protocol for generating a wide range of nG states based on the technique of weak-value amplification (WVA) (Aharonov et al., 1988; Dressel et al., 2014), without the need for explicit photon addition or subtraction operations. Our scheme relies on a weak cross-Kerr interaction between two optical modes, the signal and idler, mediated by third-order nonlinear susceptibility $\chi^{(3)}$. A nG state is conditionally generated in the signal output port from a given input signal state, heralded by single-photon detection in the idler output mode. The key element of the protocol is a postselected weak measurement of the photon number of the idler beam, which effectively induces single-photon nonlinearity via the WVA.
@@
Although the protocol features a low probability of success, it enables the high-fidelity generation of various nG states. By appropriately tuning the weak value of the photon number operator in the idler mode within experimentally feasible parameters, our protocol can generate photon-added states and displaced number states from coherent input states. In the case of a squeezed vacuum (SV) input, the scheme yields high-fidelity two-photon-added SV states, squeezed number states, and equivalent squeezed even Schrödinger cat (SC) states. When applied to SC input states, the protocol serves two distinct purposes: (i) it significantly enhances the non-Gaussianity of the initial state, especially for small-amplitude even SC states, and (ii) it enables the expansion of small SC states into large-amplitude ($|\alpha|\geq 2$) SC states while preserving high fidelity ($F>0.99$). Additionally, our approach supports the generation of a continuous family of nonclassical states with intermediate properties corresponding to different classes of input states.
The remainder of this paper is organized as follows. In Sec. II, we present our nG state generation protocol and briefly discuss its feasibility for implementation in an optical laboratory. In Secs. III and IV, we provide a detailed theoretical analysis for generating the representative nG states. We demonstrate that for coherent and SV input states, our protocol can produce high-fidelity nonclassical states, including the SPAC state, displaced single-photon state, two-photon-added squeezed vacuum (TPASV) state, squeezed photon state, and squeezed even SC state. These states are of particular interest because of their potential applications in quantum information processing. For SC input states, we show that the protocol not only enhances the non-Gaussianity of the initial state, but also enables the preparation of large-amplitude SC states with $\alpha\geq 2$ and fidelity exceeding 0.99. Finally, we conclude the paper in Sec. V. Throughout this study, we adopted natural units with $\hbar=1$. All the numerical simulations were performed using the Python package QuTiP (Johansson et al., 2012, 2013).
II Model setup
@@
A previous study (Hu et al., 2022) introduced a quantum state engineering scheme based on postselected weak measurements implemented in a medium characterized by the second-order nonlinear susceptibility of a BBO crystal. The associated interaction Hamiltonian is given by $H_{I}=\xi(a^{\dagger}b^{\dagger}-ab)$, where $a$ ($a^{\dagger}$) and $b$ ($b^{\dagger}$) are the annihilation (creation) operators of the signal and idler modes, respectively. In that proposal, the interaction strength $\chi^{(2)}$ is proportional to the second-order nonlinear susceptibility of the medium, that is $\xi\propto\chi^{(2)}$, and the signal and idle modes play the roles of the pointer and the measured system, respectively. In contrast, in many quantum optical protocols involving light-matter interactions, a different von Neumann-type interaction Hamiltonian is often employed, namely, the cross-Kerr interaction, described by
$H_{int}=\chi^{(3)}n_{a}n_{b},$ (1)
@@
where $n_{a}=a^{\dagger}a$ and $n_{b}=b^{\dagger}b$ are the number operators for the signal and idler modes, respectively, and $\chi^{(3)}$ denotes the third-order nonlinear susceptibility of the medium. As demonstrated in previous work (Feizpour et al., 2011; Hallaji et al., 2017; Matsuoka et al., 2017), if the signal and idler beams are identified as the measured system and the pointer, respectively, the interaction Hamiltonian $H_{int}$ can be used within a von Neumann-type measurement framework. Analogously to the standard weak measurement scenario, in the Hamiltonian $H_{int}$, the photon number operator $n_{a}$ serves as the observable of the measured system, while $n_{b}$ acts as the pointer variable, canonically conjugate to the phase operator $\varphi$ of the radiation field. In this section, we demonstrate the utility of the interaction Hamiltonian $H_{int}$ for preparation of quantum states via a postselected weak measurement technique.
@@
Figure 1: Schematic of the nG state generation protocol. In this scheme, a given state is prepared in the input signal mode, and a single-photon state is prepared in the idler mode, both of which pass through a Mach-Zehnder interferometer (MZI). Preselection is performed by sending the single photon through an unbalanced beam splitter (BS) with a small deviation $\epsilon$ from the ideal angle $\frac{\pi}{4}$. The signal and idler modes are then coupled via a Kerr medium. A single-photon detection event at one of the idler output ports (i.e., a click at detector $D_{1}$) heralds the desired output state in the signal mode.
@@
The schematic setup of our state generation model, based on a postselected weak measurement characterized by the Hamiltonian interaction $H_{int}$, is shown in Fig. 1. As illustrated, the proposed setup consists of a Mach-Zehnder interferometer (MZI), Kerr medium, and several optical elements. In our model, the beam splitter (BS) plays a crucial role in the execution of the protocol. The action of the BS is described by the scattering matrix (Campos et al., 1989; Agarwal, 2013).
@@
where the reflectance and transmittance of the BS are given by $\sqrt{R}=\sin\theta e^{i\phi_{\rho}}$and $\sqrt{T}=\cos\theta e^{i\varphi_{\tau}}$, respectively, satisfying $R+T=1$. We assume that the signal mode is initially prepared in an arbitrary quantum state $|\phi\rangle$ and passes through the Kerr medium, while the idle mode is initialized in a single-photon state. As shown in Fig. 1, when the input to the first BS is $|1\rangle|0\rangle$ (single photon and vacuum), the resulting state after the BS is
$|\psi_{i}\rangle$ $=\cos\theta|0\rangle_{r}|1\rangle_{t}+i\sin\theta|1\rangle_{r}|0\rangle_{t}.$ (3)
assuming $\varphi_{\tau}=0$, and $\phi_{\rho}=\frac{\pi}{2}$. The interaction between the measured system and pointer occurs via the Kerr medium, described by the time-evolution operator:
[…]
where $g=\chi^{(3)}t$ is the interaction strength, with $t=L/\upsilon$ being the interaction time, $L$ is the length of the Kerr medium, and $\upsilon$ is the speed of light in the medium. Since $g$ is typically small, it suffices to expand the evolution operator to the first order as follows:
$U=\exp\left(-ig\hat{n}_{a}\otimes\hat{n}_{b}\right)\approx 1-ig\hat{n}_{a}\otimes\hat{n}_{b}$ (5)
@@
$|\Psi^{\prime}\rangle$ $=U|\psi_{i}\rangle\otimes|\phi\rangle=\exp\left(-ig\hat{n}_{a}\otimes\hat{n}_{b}\right)|\psi_{i}\rangle\otimes|\phi\rangle$
$\approx\left[1-ig\hat{n}_{a}\otimes\hat{n}_{b}\right]|\psi_{i}\rangle\otimes|\phi\rangle$ (6)
@@
[…]
Postselecting $|\psi_{f}\rangle$ from the state $|\Psi^{\prime}\rangle$, yields the (unnormalized) final pointer state:
@@
is the weak value of the photon number operator $n_{a}=a^{\dagger}a$ for the idler beam. This expression assumes a slight deviation from a $50:50$ BS, with $\theta=\frac{\pi}{4}-\epsilon$, where $\epsilon\ll 1$ is real. For a sufficiently small $\epsilon$, the weak value can exceed unity, as observed in previous studies (Feizpour et al., 2011; Hallaji et al., 2017). The success probability of the postselection (i.e., a click at $D1$ and no click at $D2$ ) is given by $|\langle\psi_{f}|\psi_{i}\rangle|^{2}=|\epsilon|^{2}$. The normalized output state of the signal mode is
@@
where $p_{f}=|\langle\Phi|\Phi\rangle|^{2}$ is the probability of success of the post-selection. From the form of $|\Phi_{f}\rangle$, it is evident that the output state is a superposition of the initial state $|\phi\rangle$ and the state $b^{\dagger}b|\phi\rangle$. The interference between these components enhances the quality of the generated state and can lead to non-trivial quantum effects. In particular, the term $b^{\dagger}b|\phi\rangle$ corresponds to a photon subtraction followed by photon addition, a process previously identified as a powerful tool for quantum state engineering (Parigi et al., 2007). Therefore, when the coefficient $\frac{g}{2}\left(1+\frac{1}{\epsilon}\right)\gg 1$, the $b^{\dagger}b|\phi\rangle$ component dominates the output, effectively transforming the initial state $|\phi\rangle$ into a distinct quantum state with nontrivial properties and nG characteristics.
@@
As investigated in a recent WVA experiment (Li et al., 2018), the minimal rotation angle $\epsilon$ of the beam splitter (BS) can be as small as $0.52$ mrad ($5.2\times 10^{-4}$ rad). In a more recent study (Li et al., 2025), the BS was also considered to have deviated slightly from the ideal configuration $\pi/4$, with the deviation characterized by $\epsilon\propto 10^{-3}$. Generally, the non-linearity coefficient of the cross-Kerr $g$ is small (Dong et al., 2008). However, recent studies have proposed methods to engineer giant cross-Kerr nonlinearities (Wang et al., 2001; Qian et al., 2016; Kounalakis et al., 2018; van Doai et al., 2019; Doai, 2019; Mu et al., 2021; Huang et al., 2024), with a reported maximum value of $g=0.35$ (Hoi et al., 2013). Reference (Du et al., 2024) explored the implementation of various quantum logic gates based on enhanced cross-Kerr nonlinearity. These findings suggest that the weak value $\langle n_{a}\rangle_{w}$ can be significantly amplified by WVA (Feizpour et al., 2011; Hallaji et al., 2017), making it feasible to achieve the condition $\frac{g}{2}\left(1+\frac{1}{\epsilon}\right)\gg 1$ with appropriate control techniques. This result indicates that the preparation of a new family of quantum states using our proposed scheme is experimentally accessible. In the next section, we present examples that illustrate the preparation of nG states and the enhancement of non-Gaussianity for different input signal states $|\phi\rangle$.
III Non-Gaussian sates generation
III.1 photon-added coherent sate
@@ III Non-Gaussian sates generation
According to Eq. (8), if the input signal state is taken to be a coherent state, i.e., $|\phi\rangle=|\beta\rangle=D\left(\beta\right)|0\rangle$, where the displacement operator is defined as $D(\beta)=\exp\left(\beta b^{\dagger}-\beta^{\ast}b\right)$, then the normalized final state of the signal beam, after the postselected weak measurement implemented via our protocol, becomes
@@
$|\Phi^{\prime}\rangle$ $=\mathcal{N}\left[|\beta\rangle-ig\langle n_{a}\rangle_{w}b^{\dagger}b|\beta\rangle\right]=\mathcal{N}\left[|\beta\rangle-\kappa|\phi_{1}\rangle\right],$ (11)
@@
where $\kappa=i\beta g\langle n_{a}\rangle_{w}$, $|\phi_{1}\rangle=b^{\dagger}|\beta\rangle$, and the normalization constant $\mathcal{N}$ is given by
$\mathcal{N}$ $=\left[1+|\kappa|^{2}(1+|\beta|^{2})-2Re[\kappa^{\ast}\beta]\right]^{-\frac{1}{2}}.$ (12)
@@
Here, $Re[\mathcal{C}]$ denotes the real part of the complex number $\mathcal{C}$. Furthermore, using the commutation relation $[b^{\dagger},D(\beta)]=\beta^{\ast}D(\beta)$, the output signal state $|\Phi^{\prime}\rangle$ can be rewritten as
$|\Phi^{\prime}\rangle=\mathcal{N}\left[\left(1-\kappa\beta^{\ast}\right)|\beta\rangle-\kappa|\beta,1\rangle\right],$ (13)

### v2

@@ ABSTRACT
We introduce a protocol for generating non-Gaussian (nG) states via postselected weak measurement. The scheme involves injecting an arbitrary quantum state and a single photon into the signal and idler ports of an interferometer with a third-order nonlinear medium. An nG state is conditionally produced at the signal output, heralded by single-photon detection in an idler output channel. The protocol exploits a weak cross-Kerr interaction, with effective single-photon nonlinearity enhanced by weak-value amplification. By tuning the weak value of the idler photon number operator within experimentally feasible parameters, diverse nG states can be generated with high fidelity. Specific examples include photon-added coherent states, displaced and squeezed number states, and intermediate nG states from coherent and squeezed vacuum inputs. Furthermore, the protocol enables enhancement of non-Gaussianity and enlargement of Schrödinger cat (SC) states when ideal SC states are used as input. Our results provide an alternative route for conditional generation of tunable nG states, with potential applications in quantum information processing and state engineering.
INTRODUCTION
@@
Non-Gaussian (nG) states (PRXQuantum.2.030204; PRXQuantum.1.020305), defined as quantum states in the phase space that deviate from a Gaussian distribution, are often characterized by Wigner function negativity and have garnered significant interest due to their enhanced utility across various domains of quantum physics. In particular, they play a central role in continuous-variable (CV) quantum information processing (RevModPhys.77.513), including quantum computation (PhysRevLett.82.1784; PhysRevA.64.012310; PhysRevLett.97.110501; PhysRevA.93.022301; PhysRevLett.109.230503), quantum metrology (PhysRevLett.104.103602; PhysRevA.78.063828), and foundational studies in quantum theory (PhysRevA.67.012106; PhysRevA.75.052105). The generation of nG states requires either nG initial states or nG measurements; the latter alone is sufficient to produce non-Gaussianity. Among the existing techniques, nG unitary transformations and conditional measurements are the two most common approaches (PRXQuantum.2.030204). Over the past decades, considerable efforts have focused on preparing nG states using conditional measurements involving nG quantum operations, such as photon addition and/or subtraction applied to Gaussian states in optical systems (PhysRevA.82.063833; 2004; Wakui:07; PhysRevLett.97.083604; PhysRevLett.92.153601; PhysRevA.110.023703).
@@
Squeezed single-photon and single-photon-added coherent (SPAC) states are among the most representative nG states and play irreplaceable roles in CV quantum information processing. These states can be prepared via nG operations such as single-photon addition and subtraction, applied to coherent and squeezed vacuum (SV) states, respectively. Another paradigmatic example is the Schrödinger cat (SC) state (2006) inspired by Schrödinger’s famous thought experiment (schrodinger1935gegenwartige). Current methods for generating SC states include photon-number-resolving detection (PhysRevA.55.3184; PhysRevA.59.1658) and weak and strong nonlinear optical interactions (PhysRevLett.57.13; PhysRevA.59.4095). Large-amplitude SC states characterized by coherent amplitudes $|\alpha|\geq 2$, where $\alpha$ is the complex amplitude, are particularly valuable for probing quantum foundations and enabling quantum information protocols (PhysRevA.64.022313; PhysRevA.68.022321; PhysRevA.69.022315). However, SC states generated via photon subtraction or third-order optical nonlinearities typically have small amplitudes and fall short of the requirements imposed by advanced quantum tasks.
@@
From a practical perspective, the purity of a state is critical for effective implementation of many quantum information protocols. However, due to limitations in photon detection efficiency and photon-number-resolving capabilities, generating high-fidelity nG states remains significantly more challenging than producing their Gaussian counterparts. This difficulty underscores the limitations of existing methods for preparing high-quality nG states. To address this challenge, one can either enhance the non-Gaussianity of a given state through optimization techniques or develop alternative approaches capable of generating nG states with high purity and fidelity tailored for specific applications.
@@
In this study, we propose an alternative protocol for generating a wide range of nG states based on the technique of weak-value amplification (WVA) (PhysRevLett.60.1351; RevModPhys.86.307), without the need for explicit photon addition or subtraction operations. Our scheme relies on a weak cross-Kerr interaction between two optical modes, the signal and idler, mediated by third-order nonlinear susceptibility $\chi^{(3)}$. A nG state is conditionally generated in the signal output port from a given input signal state, heralded by single-photon detection in the idler output mode. The key element of the protocol is a postselected weak measurement (WM) of the photon number of the idler beam, which effectively induces single-photon nonlinearity via the WVA.
@@
Although the protocol features a low probability of success, it enables the high-fidelity generation of various nG states. By appropriately tuning the weak value of the photon number operator in the idler mode within experimentally feasible parameters, our protocol can generate photon-added states and displaced number states from coherent input states. In the case of a SV input, the scheme yields high-fidelity two-photon-added squeezed vacuum (TPASV) states, squeezed number states, and equivalent squeezed even SC states. When applied to SC input states, the protocol serves two distinct purposes: (i) it significantly enhances the non-Gaussianity of the initial state, especially for small-amplitude even SC states, and (ii) it enables the expansion of small SC states into large-amplitude ($|\alpha|\geq 2$) SC states while preserving high fidelity ($F>0.99$). Additionally, our approach supports the generation of a continuous family of nonclassical states with intermediate properties corresponding to different classes of input states.
RESULTS
A Versatile Protocol via Weak Measurement
@@
A previous study (PhysRevA.105.022608) introduced a quantum state engineering scheme based on postselected WMs implemented in a medium characterized by the second-order nonlinear susceptibility of a BBO crystal. The associated interaction Hamiltonian is given by $H_{I}=\xi(a^{\dagger}b^{\dagger}-ab)$, where $a$ ($a^{\dagger}$) and $b$ ($b^{\dagger}$) are the annihilation (creation) operators of the signal and idler modes, respectively. In that proposal, the interaction strength $\chi^{(2)}$ is proportional to the second-order nonlinear susceptibility of the medium, that is $\xi\propto\chi^{(2)}$, and the signal and idler modes play the roles of the pointer and the measured system, respectively. In contrast, in many quantum optical protocols involving light-matter interactions, a different von Neumann-type interaction Hamiltonian is often employed, namely, the cross-Kerr interaction, described by
$H_{int}=\chi^{(3)}n_{a}\otimes n_{b},$ (1)
@@
where $n_{a}=a^{\dagger}a$ and $n_{b}=b^{\dagger}b$ are the number operators for the signal and idler modes, respectively, and $\chi^{(3)}$ denotes the third-order nonlinear susceptibility of the medium. As demonstrated in previous works (PhysRevLett.107.133603; Nature2017; RN11), if the signal and idler beams are identified as the measured system and the pointer, respectively, the interaction Hamiltonian $H_{int}$ can be used within a von Neumann-type measurement framework. Analogously to the standard WM scenario, in the Hamiltonian $H_{int}$, the photon number operator $n_{a}$ serves as the observable of the measured system, while $n_{b}$ acts as the pointer variable, canonically conjugate to the phase operator $\varphi$ of the radiation field. In this section, we demonstrate the utility of the interaction Hamiltonian $H_{int}$ for preparation of quantum states via a postselected WM technique.
@@
Figure 1: Schematic of the nG state generation protocol. In this scheme, a given state is prepared in the input signal mode, and a single-photon state is prepared in the idler mode, both of which pass through a MZI. Preselection is performed by sending the single photon through an unbalanced BS with a small deviation $\epsilon$ from the ideal angle $\pi/4$. The signal and idler modes are then coupled via a Kerr medium. A single-photon detection event at one of the idler output ports (i.e., a click at detector $D_{1}$) heralds the desired output state in the output signal mode.
@@
The schematic setup of our state generation model, based on a postselected WM characterized by the Hamiltonian interaction $H_{int}$, is shown in Fig. 1. As illustrated, the proposed setup consists of a Mach-Zehnder interferometer (MZI), Kerr medium, and several optical elements. In our model, the beam splitter (BS) plays a crucial role in the execution of the protocol. The action of a two-mode BS is described by the scattering matrix $U_{BS}$ (PhysRevA.40.1371; Agarwal2013) with
@@
where the reflectance and transmittance of the BS are given by $\sqrt{R}=\sin\theta$ and $\sqrt{T}=\cos\theta$, respectively, satisfying $R+T=1$. We assume that the signal mode is initially prepared in an arbitrary quantum state $|\phi\rangle$ and passes through the Kerr medium, while the idler mode is initialized in a single-photon state. As shown in Fig. 1, when the input to the first BS is $|1\rangle_{1}|0\rangle_{2}$, the resulting state after the BS is
$|\psi_{i}\rangle$ $=\cos\theta_{1}|0\rangle_{r}|1\rangle_{t}+i\sin\theta_{1}|1\rangle_{r}|0\rangle_{t},$ (3)
assuming $\varphi_{\tau}=0$, and $\phi_{\rho}=\pi/2$. The subindices $r$ and $t$ of the states represent the existence of single-photon in reflected and transmitted paths of the interferometer after the first BS. The interaction between the measured system and pointer occurs via the Kerr medium, described by the time-evolution operator:
$U$ $=\exp\left(-\frac{i}{\hbar}\int_{0}^{t}H_{I}d\tau\right)=\exp\left(-ign_{a}\otimes n_{b}\right),$ (4)
where $g=\chi^{(3)}t$ is the interaction strength, with $t=L/\upsilon$ being the interaction time, where $L$ is the length of the Kerr medium, and $\upsilon$ is the speed of light in the medium. If $g$ is typically small and mean photon numbers in $a$ and $b$ modes both not large, it suffices to expand the evolution operator to the first order as follows:
$U=\exp\left(-ign_{a}\otimes n_{b}\right)\approx 1-ign_{a}\otimes n_{b}.$ (5)
@@
$|\Psi^{\prime}\rangle$ $=U|\psi_{i}\rangle\otimes|\phi\rangle=\exp\left(-ign_{a}\otimes n_{b}\right)|\psi_{i}\rangle\otimes|\phi\rangle$
$\approx\left[1-ign_{a}\otimes n_{b}\right]|\psi_{i}\rangle\otimes|\phi\rangle.$ (6)
Since the idler mode in our scheme contains only one photon, the validity of this approximation requires $g\langle n_{b}\rangle\ll 1$ and $\frac{1}{2}g^{2}\langle n_{b}^{2}\rangle\ll 1$. These conditions can be satisfied with weak nonlinearities and input states of moderate photon numbers. Here, $\langle n_{b}\rangle$ and $\langle n_{b}^{2}\rangle$ represent the expectation values of the operators $n_{b}$ and $n_{b}^{2}$ for the initial signal state $|\phi\rangle$, respectively.
@@
$|\psi_{f}\rangle$ $=\text{$\frac{1}{\sqrt{2}}(|0\rangle_{r}|1\rangle_{t}-i|1\rangle_{r}|0\rangle_{t})$,}$ (7)
where we assume $\varphi_{\tau}=0$, $\phi_{\rho}=\pi/2$ and $\theta_{2}=\pi/4$ for a balanced BS. The details of deriviations of pre- and pre-selected states are provided in Methods. Postselecting $|\psi_{f}\rangle$ from the state $|\Psi^{\prime}\rangle$, yields the (unnormalized) final pointer state:
@@
is the weak value of the photon number operator $n_{a}=a^{\dagger}a$ for the idler beam. This expression assumes a slight deviation from a $50:50$ BS, with $\theta_{1}=\pi/4-\epsilon$, where $\epsilon\ll 1$ is real. For a sufficiently small $\epsilon$, the weak value can exceed unity, as observed in previous studies (PhysRevLett.107.133603; Nature2017). The success probability of the postselection (i.e., a click at $D1$ and no click at $D2$ ) is given by $|\langle\psi_{f}|\psi_{i}\rangle|^{2}=\epsilon^{2}$. This indicates that whenever the postselection succeeds our protocol amplifies the mean photon number of the idler mode as if there exist more than single photon inside the interferometer. Consequently, the signal beam experience a cross-phase shift equivalent to that of many photons.
In the above postselection process we assume that the second BS is ideally $50:50$ balanced. However, in practical implementation processes of our scheme it is necessary to take into account the imperfections of related optical elements. If both BSs are slightly unbalanced—for example, with their transmission-to-reflection ratios characterized by the parameters $\theta_{1}=\pi/4-\epsilon_{1}^{\prime}$ and $\theta_{2}=\pi/4\pm\epsilon_{2}^{\prime},$ respectively–the weak value obtained in above case changed to $1/2+1/2(\epsilon_{1}^{\prime}\mp\epsilon_{2}^{\prime})$, where $\epsilon_{1}^{\prime},\epsilon_{2}^{\prime}\ll 1$. Even under such imbalanced conditions, by tuning the system parameters to satisfy the phase-matching condition $\epsilon_{1}^{\prime}\mp\epsilon_{2}^{\prime}=\epsilon$, one can still obtain output results identical to those of an ideal 50 : 50 BS. This demonstrates the robustness of our protocol to such imperfections, as they can be compensated for by careful calibration of the optical apparatus.
After the above postselected WM procedure, the normalized output state of the signal mode is
@@
where the overall success probability $p_{f}$ of the protocol for generating the signal output state with single-trial is given by the general form $p_{f}=\epsilon^{2}|\mathcal{N}|^{-2}.$ Here, $\epsilon$ is the small deviation angle of the first BS from $\pi/4$, and $\mathcal{N}$ is the normalization constant of the final state $|\Phi\rangle$, which depends on the specific input state $|\phi\rangle$ and system parameters ($g,\epsilon$). A key strength of our approach is that the success probability across all types of input states follows a universal form, allowing for a unified feasibility analysis.
@@
From the form of $|\Phi_{f}\rangle$, it is evident that the output state is a superposition of the initial state $|\phi\rangle$ and the state $b^{\dagger}b|\phi\rangle$. The interference between these components enhances the quality of the generated state and can lead to nontrivial quantum effects. In particular, the term $b^{\dagger}b|\phi\rangle$ corresponds to a photon subtraction followed by photon addition, a process previously identified as a powerful tool for quantum state engineering (2007PV). Therefore, when the coefficient $\frac{g}{2}\left(1+\frac{1}{\epsilon}\right)\gg 1$, the $b^{\dagger}b|\phi\rangle$ component dominates the output, effectively transforming the initial state $|\phi\rangle$ into a distinct quantum state with nontrivial properties and nG characteristics.
@@
As investigated in a recent WVA experiment (PhysRevA.97.033851), the minimal rotation angle $\epsilon$ of the BS can be as small as $0.52$ mrad ($5.2\times 10^{-4}$ rad). In a more recent study (PhysRevA.111.042425), the BS was also considered to have deviated slightly from the ideal configuration $\pi/4$, with the deviation characterized by $\epsilon\propto 10^{-3}$. Generally, the non-linearity coefficient of the cross-Kerr $g$ is small (DONG20085677). However, recent studies have proposed methods to engineer giant cross-Kerr nonlinearities (PhysRevLett.87.073601; 2016; 2018; vanDoai:19; Doai_2019; PhysRevA.103.043709; 2024), with a reported maximum value of $g=0.35$ (PhysRevLett.111.053601). Reference (Du2024kerreffectbased) explored the implementation of various quantum logic gates based on enhanced cross-Kerr nonlinearity. These findings suggest that the weak value $\langle n_{a}\rangle_{w}$ can be significantly amplified by WVA (PhysRevLett.107.133603; Nature2017), making it feasible to achieve the condition $\frac{g}{2}\left(1+\frac{1}{\epsilon}\right)\gg 1$ with appropriate control techniques. As described, in our protocol, the average photon number $\langle n_{b}\rangle$ is determined by the input state $|\phi\rangle$. The parameters of this input state can be adjusted to enhance the weight of the $n_{b}$ term in the post-measurement state $|\Phi_{f}\rangle$ via WVA, while maintaining the weak coupling condition $g\langle n_{b}\rangle\ll 1$. Our robust parameter analysis thereby confirms the preparation of a new family of quantum states with our proposed scheme is experimentally feasible.
Below, we present examples that illustrate the preparation of nG states and the enhancement of non-Gaussianity for different input signal states $|\phi\rangle$.
Non-Gaussian States Generation
@@
As a first example, we demonstrate the preparation of a state closely resembling a photon-added coherent state. According to Eq. (8), if the input signal state is taken to be a coherent state, i.e., $|\phi\rangle=|\beta\rangle=D\left(\beta\right)|0\rangle$, where the displacement operator is defined as $D(\beta)=\exp\left(\beta b^{\dagger}-\beta^{\ast}b\right)$, then the normalized final state of the signal beam, after the postselected WM implemented via our protocol, becomes
@@
$|\Phi^{\prime}\rangle$ $=\mathcal{N}_{c}\left[|\beta\rangle-ig\langle n_{a}\rangle_{w}b^{\dagger}b|\beta\rangle\right]=\mathcal{N}_{c}\left[|\beta\rangle-\kappa|\phi_{1}\rangle\right],$ (11)
@@
where $\kappa=i\beta g\langle n_{a}\rangle_{w}$, $|\phi_{1}\rangle=b^{\dagger}|\beta\rangle$, and the normalization constant $\mathcal{N}_{c}$ is given by
$\mathcal{N}_{c}$ $=\left[1+|\kappa|^{2}(1+|\beta|^{2})-2Re[\kappa^{\ast}\beta]\right]^{-\frac{1}{2}}.$ (12)
@@
Here, $Re[\mathcal{C}]$ denotes the real part of the complex number $\mathcal{C}$. Furthermore, using the commutation relation $b^{\dagger},D(\beta)]=\beta^{\ast}D(\beta)$, the output signal state $|\Phi^{\prime}\rangle$ can be rewritten as
$|\Phi^{\prime}\rangle=\mathcal{N}_{c}\left[\left(1-\kappa\beta^{\ast}\right)|\beta\rangle-\kappa|\beta,1\rangle\right],$ (13)